OpenStudy (anonymous):

given f(x)=8x+13 and g(x)=x^2-5x find the function (f+g)(x).

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x^2 + 3x +13

OpenStudy (anonymous):

just add the two

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x)=x^2-5x+8x+13=x^2+3x+13... did u understand samephone?

OpenStudy (anonymous):

-13 at the end?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no there's no -13 , its because ur just adding them :)

OpenStudy (anonymous):

so its x^2-3X13?

OpenStudy (anonymous):

thx soooo much i understand!